"What is the height of a building with 6 floors given by the formula h= 20.6 + 3.5n"
Thank you.

Answer by mathchemprofessor(65)

The height of a building with 6 floors given by the formula h= 20.6 + 3.5n
When n=6,
the height of a building with 6 floors given by the formula h= 20.6 + 3.5*6
=20.6+21
=41.6